# MalAlgoQA: A Pedagogical Approach for Evaluating Counterfactual Reasoning Abilities

### v1

@@ Abstract
This paper introduces MalAlgoQA, a novel dataset designed to evaluate the counterfactual reasoning capabilities of Large Language Models (LLMs) through a pedagogical approach. The dataset comprises mathematics and reading comprehension questions, each accompanied by four answer choices and their corresponding rationales. We focus on the incorrect answer rationales, termed "malgorithms", which highlights flawed reasoning steps leading to incorrect answers and offers valuable insights into erroneous thought processes. We also propose the Malgorithm Identification task, where LLMs are assessed based on their ability to identify corresponding malgorithm given an incorrect answer choice. To evaluate the model performance, we introduce two metrics: Algorithm Identification Accuracy (AIA) for correct answer rationale identification, and Malgorithm Identification Accuracy (MIA) for incorrect answer rationale identification. The task is challenging since state-of-the-art LLMs exhibit significant drops in MIA as compared to AIA. Moreover, we find that the chain-of-thought prompting technique not only fails to consistently enhance MIA, but can also lead to underperformance compared to simple prompting. These findings hold significant implications for the development of more cognitively-inspired LLMs to improve their counterfactual reasoning abilities, particularly through a pedagogical perspective where understanding and rectifying student misconceptions are crucial. The data and code can be found at https://github.com/luffycodes/MalAlgoQA-Dataset.
@@ 1 Introduction
Counterfactual reasoning refers to the ability to consider alternative scenarios and infer the cause of an event based on hypothetical situations. It is a fundamental aspect of human cognition that plays a crucial role in problem-solving, decision-making, and learning [5, 13, 14]. Although large language models (LLMs) have demonstrated remarkable performance in various natural language processing tasks, their ability to engage in counterfactual reasoning remains relatively underexplored. Counterfactual reasoning requires a degree of hypothetical and abstract thinking, which poses a significant challenge for current LLMs, as they are generally trained to predict outcomes based on facts, rather than imagining alternative scenarios that might not be explicitly present in the training data [3, 9, 19].
@@
To fill this gap, we introduce MalAlgoQA, a novel dataset designed to evaluate the counterfactual reasoning capabilities of LLMs through a pedagogical approach with multiple-choice question answering. MalAlgoQA consists of a diverse set of 807 mathematics and 290 reading comprehension questions spanning various grade levels (3-11), content classifications (e.g., algebra, geometry, number & operations), and Depth of Knowledge levels (DOK) (1-3) [25]. Each question is accompanied by a set of answer choices and their associated rationales, which we term "malgorithms" (inspired by the concept of "mal-rules" proposed by [11]). Malgorithms represent the flawed reasoning that leads to an incorrect answer, revealing the underlying thought process that causes a student to arrive at the incorrect conclusion.
With MalAlgoQA, we propose the Malgorithm Identification task, which involves identifying the rationale behind the selection of a particular answer choice. When the given answer choice is correct, a model needs to employ causal reasoning to determine the rationale or algorithm that justified the correct selection. On the other hand, when the given answer choice is incorrect, the task demands counterfactual reasoning, as the model must identify the flawed reasoning or malgorithm that led to the selection of an incorrect option. In this case, the model must consider hypothetical scenarios and reason about the cause of an incorrect choice, which may not align with the facts. To illustrate the Malgorithm Identification task, consider a simple question: "What is 1 + 2 × 3 + 4?" with choices A. 11, B. 13, and C. 21. The rationales are: "Worked left to right" (corresponding to choice B), "Applied PEASMD rule instead of PEMDAS" (corresponding to choice C), and "Applied PEMDAS rule" (corresponding to the correct choice A). Given an answer choice, the model should identify the rationale explaining that choice.
@@
In order to evaluate model performance, we introduce two key metrics: Algorithm Identification Accuracy (AIA), which measures accuracy in identifying the rationale behind a given correct answer. Malgorithm Identification Accuracy (MIA), of particular interest, evaluates the model’s counterfactual causal reasoning abilities by quantifying accuracy in identifying the rationale behind an incorrect answer. Our experiments reveal that GPT-4o [10], a state-of-the-art LLM, achieves an AIA of $95.7\%$, demonstrating its proficiency in this task. However, when it comes to MIA, GPT-4o’s performance drops significantly, with an accuracy of only $66.1\%$. This pattern, where accuracy is higher for correct answer rationales and lower for incorrect answer rationales, is consistent across other language models, such as LLaMA-3-70B [1] and GPT-3.5 [18]. It emphasizes the challenge of counterfactual reasoning for LLMs, especially on finding the underlying rationale behind incorrect answer choices.
Our experiments also reveal an unexpected result that Chain-of-Thought (CoT) prompting not only fails to consistently enhance MIA, it sometimes even underperforms compared to simple prompting. For example, GPT-4o achieves an MIA of 68.6% with CoT prompting compared to 82.8% with simple prompting for Reading. These results suggest that the reasoning steps provided by CoT prompting can be detrimental for robust counterfactual reasoning.
Our experiments reveal a decline in LLM performance with increasing DOK and grade level, highlighting challenges in handling complex questions and deep reasoning for the Malgorithm identification task. Probability emerges as the most challenging content area, while geometry is the least. This trend underscores the struggle of LLMs to sustain high performance as cognitive demands escalate.
Additionally, our experiments indicate that LLMs performance declines as the DOK and grade level increases, suggesting that the complexity of the questions and the depth of reasoning pose greater challenges for the Malgorithm identification task. Among the content areas, Probability proves to be the most difficult for LLMs, with Geometry being the least. This pattern shows that as the cognitive demands of the tasks increase, LLMs struggle more to maintain high performance levels. Overall, our findings further emphasize the need for research and development of models that can effectively engage in counterfactual reasoning.
2 Dataset: MalAlgoQA
@@
We introduce MalAlgoQA, a real-world dataset consisting of multiple-choice questions in both mathematics and reading comprehension for students in grade 3 through 11. The dataset is designed to evaluate the counterfactual reasoning abilities of LLMs through the task of analyzing the rationales and reasons behind incorrect answer choices, a process we refer to as “Malgorithm identification". The MalAlgoQA dataset provides valuable resources to assess the counterfactual reasoning capabilities of LLMs through a pedagogical approach.
@@ 2.1 Dataset Overview
MalAlgoQA comprises two multiple-choice question sets: a mathematics and a reading comprehension question set. The mathematics question set contains 807 questions, while the reading comprehension question set has 290 questions. Each question is associated with four answer choices, as well as their corresponding rationales. Reading question set also provides a passage for each question. Table 1 provides an example of math question, with more examples available in BC. Additionally, the dataset provides rich metadata including the grade level, content classification, and Depth of Knowledge (DOK) level for each question.
@@ 2.1.3 Depth of Knowledge (DOK) Levels
Each question in MalAlgoQA is annotated with a DOK level, which indicates the cognitive complexity of the question. The DOK levels are derived from Norman L. Webb’s taxonomy [25] and range from I to III, with higher levels representing increased cognitive complexity. Table 4 shows the distribution of questions across DOK levels for both question sets.
@@ 3 Methodology
In this section, we present the task formulation for Malgorithm Identification, which aims to evaluate the performance of LLMs by identifying the relationship between answer choices and their underlying rationales.
3.1 Definitions and Task Formulation
@@
Answer Rationales:
{formatted_rationales}
@@
Table 5: Prompt for the Malgorithm Identification experiment, where the LLM identifies the relationship between answer choices and rationales in a multiple-choice question-answering setting. The task involves identifying a "malgorithm", a rationale that represents flawed reasoning leading to an incorrect answer choice. The prompt provides the question, an answer choice (either correct or incorrect), and all rationales, asking the LLM to identify the rationale corresponding to the given choice. The highlighted sentence instructs the model to provide step-by-step reasoning, an instruction added for the Chain-of-Thought prompting strategy.
@@
We introduce the term "malgorithm" to describe a rationale that contains a combination of correct reasoning steps and one or more flawed steps, which we call "mal-rules,". These "mal-rules" lead to incorrect answer choices. The term "mal-rule" is derived from the cognitive error literature [4, 23, 11], which refers to the flawed or incorrect rules that students may apply when solving problems, resulting in systematic errors. In our context, a malgorithm is a reasoning chain that consists of both correct rules and mal-rules, with the presence of mal-rules ultimately leading to the selection of an incorrect answer choice.
Formally, let $R_{i}^{correct}\subseteq R_{i}$ be the set of rationales that correspond to correct answer choices for question $q_{i}$, and let $R_{i}^{incorrect}\subseteq R_{i}$ be the set of rationales that correspond to incorrect answer choices. A rationale $r_{ij}\in R_{i}^{incorrect}$ is considered a malgorithm if it contains a combination of correct reasoning steps and one or more mal-rules that lead to the selection of an incorrect answer choice.
@@ 3.2 Evaluation Metrics
1. Algorithm Identification Accuracy (AIA): This metric measures the model’s ability to identify the rationale or algorithm that aligns with a correct answer choice. Essentially, AIA evaluates how well the model recognizes and follows a chain of reasoning steps that lead to the correct answer choice.
2. Malgorithm Identification Accuracy (MIA): This metric measures the model’s ability to identify the flawed reasoning or malgorithm that aligns with an incorrect answer choice. Specifically, MIA evaluates the model’s capacity to identify faulty reasoning chains that contain a combination of correct steps and one or more mal-rules, which ultimately lead to a wrong answer.
These metrics are crucial for analyzing reasoning abilities the LLMs. AIA is essential for understanding the models’ thought processes behind correct problem-solving strategies. In contrast, MIA evaluates the model’s ability identifying flawed reasoning and counterfactual logic that lead to incorrect answers. The combination of AIA and MIA enables us to analyze the models’ performance in both causal and counterfactual reasoning.
4 Experiments
@@
Table 6: Performance results of different Large Language Models (LLMs) on the Malgorithm Identification task, presented separately for Reading and Math subjects. The results are organized according to the type of prompting (Simple or CoT), and the number of shots (0 or 5). The metrics evaluated include Multiple-Choice Question (MCQ) accuracy, Algorithm Identification Accuracy (AIA), and Malgorithm Identification Accuracy (MIA). AIA measures the model’s ability to identify the correct rationale or algorithm that aligns with the given correct answer choice, whereas MIA measures the model’s ability to identify the flawed reasoning or malgorithm that aligns with the given incorrect answer choice. It can be observed that achieving higher scores in MIA is generally more challenging than AIA, highlighting the complexity of counterfactual reasoning tasks.
@@ 4.1 Experimental Setup
We conducted experiments on the Malgorithm Identification task with four state-of-the-art LLMs: GPT-4o, GPT-3.5, LLaMA3-70B and LLaMA3-8B. [^1] The specific models we use are GPT-4o: gpt-4o-2024-05-13, GPT-3.5: gpt-3.5-turbo-0125, LLaMA3-70B: Meta-LLaMA3-70B-Instruct, LLaMA3-8B: Meta-LLaMA3-8B-Instruct To evaluate the LLMs performance, we use AIA and MIA metrics (described in section 3.2).
@@
For each model, we evaluated with four experimental settings: (1) Simple prompting, where the LLMs are prompted to provide only the final answer; (2) Chain-of-Thought (CoT) prompting, where the LLMs are prompted to provide step-by-step reasoning and show the intermediate steps before giving the final answer, as shown in Table 5; (3) Zero-shot learning, where the LLMs are provided without any additional examples; (4) Few-shot learning, where the LLMs are provided with a small number of examples demonstrating the desired reasoning process, in addition to the standard prompts. Due to context length constraints, we conducted few-shot experiments only on the Math dataset.
4.2 Effectiveness of CoT Prompting
In contrast to simple prompting strategies, Chain-of-Thought prompting requires intermediate reasoning steps that guide the model towards the final answer. CoT has demonstrated success in enhancing the reasoning capabilities of language models across various tasks [27, 7]. Nevertheless, its effectiveness in identifying flawed reasoning behind incorrect answers remain unexplored. Our objective is to determine whether the additional reasoning steps provided by CoT prompting could enhance the models ability to identify malgorithms, which is crucial for assessing models’ robustness in counterfactual reasoning and understanding student mistakes.
In Table 6, we observe that for GPT-4o, Simple prompting performs similarly to CoT prompting in detecting malgorithms for Math (65.10 vs 64.99). However, in Reading dataset, Simple prompting significantly outperforms CoT prompting in identifying malgorithms (82.80 vs 68.63). In terms of AIA, CoT prompting generally performs as well or better than Simple prompting across both subjects.. For LLaMA3-70B, Simple prompting shows better MIA than CoT prompting across both Math (50.6 vs 48.31) and Reading (60.0 vs 58.4), while AIA results remain relatively comparable.
The results demonstrates that while CoT prompting could help the identification of correct algorithms, its additional reasoning steps are not necessarily beneficial for recognizing flawed reasoning when the given answer choice is incorrect. One possible explanation is that the structured nature of CoT prompting might constrain the model’s ability to explore alternative scenarios and reason about the cause of the incorrect choice. For counterfactual reasoning, the model needs to consider possibilities that deviate from the correct reasoning path, and the intermediate steps in CoT prompting may inadvertently anchor the model to a specific line of thought, making it harder to identify the flawed reasoning behind incorrect choices. Furthermore, the models’ training data may not have sufficient examples of flawed reasoning patterns, making it challenging for the models to generate appropriate intermediate reasoning steps that lead to the identification of malgorithms, even with CoT prompting. These findings underscore the importance of malgorithm identification, which evaluates LLMs on incorrect answer choices to enhance their robustness in counterfactual reasoning.
4.3 Comparison of LLMs
Table 6 shows that malgorithm detection remains a challenging task, even for GPT-4o. This difficulty indicates room for improvement in counterfactual reasoning with LLMs. Although GPT-4o achieve over 95 in answering the multiple-choice questions in MalAlgoQA, the MIA performance drops below 70. This decline suggests that the underlying issue lies in counterfactual reasoning capabilities of the models instead of the complexity of the questions.
Furthermore, larger models such as GPT-4o and LLaMA3-70B significantly outperform smaller ones in the MIA task, while in the AIA task, performance levels are comparable across model sizes. For instance, in the Math subject with zero-shot CoT prompting, GPT-4o scored 66.10 compared to 14.74 from GPT-3.5, showcasing the large disparity. In contrast, for the AIA task, GPT-4o achieved a score of 95.65 with zero-shot CoT prompting, only marginally better than GPT-3.5’s 86.28. This suggests that the AIA task is less dependent on model size. The MIA task, which requires models to delve deeper into the underlying rationale of incorrect choices, showcases the emergent abilities of large language models [26]. This indicates that counterfactual reasoning tasks are more solvable with the enhanced capabilities of larger models.
Additionally, smaller models obtain more substantial gains with few-shot learning, especially for the MIA performance. This result suggests that for the task of malgorithm identification, smaller models could benefit more from additional context.
5 Discussions

### v2

# MalAlgoQA: Pedagogical Evaluation of Counterfactual Reasoning in Large Language Models and Implications for AI in Education

@@ Abstract
This paper introduces MalAlgoQA, a novel dataset designed to evaluate the counterfactual reasoning capabilities of Large Language Models (LLMs) through a pedagogical approach. The dataset comprises mathematics and reading comprehension questions, each accompanied by four answer choices and their corresponding rationales. At the heart of MalAlgoQA are “malgorithms” - rationales behind incorrect answer choices that represent flawed yet logically coherent reasoning paths. These malgorithms serve as counterfactual scenarios, allowing us to assess an LLM’s ability to identify and analyze flawed reasoning patterns. We propose the Malgorithm Identification task, where LLMs are assessed based on their ability to identify corresponding malgorithm given an incorrect answer choice. To evaluate the model performance, we introduce two metrics: Algorithm Identification Accuracy (AIA) for correct answer rationale identification, and Malgorithm Identification Accuracy (MIA) for incorrect answer rationale identification. Our experiments reveal that state-of-the-art LLMs exhibit significant performance drops in MIA compared to AIA, highlighting the challenges in counterfactual reasoning. Surprisingly, we find that the chain-of-thought prompting technique not only fails to consistently enhance MIA but can sometimes lead to underperformance compared to simple prompting. These findings have important implications for developing LLMs with improved counterfactual reasoning, particularly relevant for AI-powered tutoring systems, where identifying and addressing student misconceptions is essential. MalAlgoQA dataset is available here.
MalAlgoQA: Pedagogical Evaluation of Counterfactual Reasoning in
Large Language Models and Implications for AI in Education
Naiming Liu Rice University nl35@rice.edu                        Shashank Sonkar Rice University ss164@rice.edu
MyCo Le Rice University msl12@rice.edu                        Richard G. Baraniuk Rice University richb@rice.edu
footnotetext: Equal contribution.
@@ 1 Introduction
Evaluating the counterfactual reasoning capabilities of Large Language Models (LLMs) – their ability to consider “what if” scenarios and infer causes based on hypothetical situations – remains a significant challenge despite their impressive performance in various natural language tasks (Pearl, 2019; Rafetseder et al., 2021; Sonkar and Baraniuk, 2023). We introduce a novel method to evaluate these capabilities, drawing inspiration from educational assessment techniques. Specifically, we leverage the practice of distractor choice generation, which involves creating plausible but incorrect answer options by envisioning hypothetical scenarios and logical yet flawed reasoning paths. This approach mimics how educators anticipate and model potential student misconceptions, engaging in counterfactual thinking about alternative cognitive processes that could lead to specific errors.
@@
Building on this foundation, we present MalAlgoQA, a dataset designed to challenge LLMs’ ability to reason about hypothetical situations and identify flawed logical pathways. MalAlgoQA consists of a diverse set of 807 mathematics and 290 reading comprehension questions spanning various grade levels (3-11), content classifications (e.g., algebra, geometry, number & operations), and Depth of Knowledge levels (DOK) (1-3) (Webb, 2002). Each question is accompanied by a set of answer choices and their associated rationales.
Central to our approach is the concept of “malgorithms” - a term we coin to describe the flawed reasoning paths that lead to incorrect answers (name is inspired by the concept of “mal-rules” proposed by Payne and Squibb (1990)). These malgorithms represent deterministic yet erroneous thought processes, mirroring the way students might logically follow a series of steps based on a misunderstanding or misapplication of concepts. We propose the Malgorithm Identification task to assess LLMs’ ability to reason about both correct and incorrect answer rationales. In this task, given a question and a specific answer choice, the model must identify the underlying rationale (algorithm or malgorithm) that led to that choice. This approach challenges LLMs to engage in both causal and counterfactual reasoning. To illustrate, consider a simple question: “What is 1 + 2 × 3 + 4?” with choices A. 11, B. 13, and C. 21. The rationales are: “Worked left to right” (corresponding to choice B), “Applied PEASMD rule instead of PEMDAS” (corresponding to choice C), and “Applied PEMDAS rule” (corresponding to the correct choice A). Given an answer choice, the model should identify the rationale explaining that choice.
@@
In order to evaluate model performance, we introduce two key metrics: Algorithm Identification Accuracy (AIA), which measures accuracy in identifying the rationale behind a given correct answer. Malgorithm Identification Accuracy (MIA), of particular interest, evaluates the model’s counterfactual causal reasoning abilities by quantifying accuracy in identifying the rationale behind an incorrect answer. Our experiments reveal that GPT-4o (OpenAI, 2023) achieves an AIA of $95.7\%$, demonstrating its proficiency in this task. However, when it comes to MIA, GPT-4o’s performance drops significantly, with an accuracy of only $66.1\%$. This pattern, where accuracy is higher for correct answer rationales and lower for incorrect answer rationales, is consistent across other language models, such as LLaMA-3-70B (Dubey et al., 2024) and GPT-3.5 (Schulman et al., 2022). It emphasizes the challenge of counterfactual reasoning for LLMs, especially on finding the underlying rationale behind incorrect answer choices. Our experiments also reveal an unexpected result that Chain-of-Thought (CoT) prompting Wei et al. (2022b); Kojima et al. (2022) not only fails to consistently enhance MIA, it sometimes even underperforms compared to simple prompting.
The implications of our findings reveal critical challenges for AI in education. The stark performance gap between AIA and MIA across all models (e.g., GPT-4o’s $95.7\%$ AIA vs. $66.1\%$ MIA) exposes fundamental limitations in LLMs’ ability to understand student misconceptions. This discrepancy suggests that current LLM tutoring systems fail to identify and address the root causes of student errors. Such limitations necessitate a rethinking of feedback mechanisms in LLM-powered educational tools. Their struggle with malgorithm identification (e.g., GPT-3.5’s dramatic drop from $86.3\%$ AIA to $14.74\%$ MIA in Math) indicates they may offer incomplete or potentially misleading guidance when addressing student mistakes. These insights have profound implications for automated grading and assessment. The significant performance drop in MIA across all models raises concerns about the reliability of AI-driven systems in evaluating open-ended questions or complex problem-solving tasks.
Our work makes two key contributions to the field of LLMs and AI for education, with critical implications for developing more effective AI-powered tutoring systems. First, we introduce MalAlgoQA in section 2, a novel dataset inspired by pedagogical distractor generation techniques, specifically designed to evaluate the counterfactual reasoning capabilities of LLMs. This approach mimics how educators create plausible but incorrect answer options, challenging LLMs to engage in the type of reasoning essential for understanding and addressing student misconceptions. Second, we propose the Malgorithm Identification task in section 3, along with two associated evaluation metrics: AIA and MIA. Our findings in sections 4 and 5 reveal a significant performance gap between these metrics across various LLMs, highlighting the challenge these models face in counterfactual reasoning. We discuss the implications of our findings for AI-powered educational tool in section 6. By developing methods to evaluate LLMs’ ability to identify and understand student misconceptions, our work paves the way for more effective AI tutoring systems that can accurately diagnose and address individual students’ conceptual gaps Liu et al. (2023); Sonkar et al. (2023, 2024c).
2 MalAlgoQA Dataset
@@
We introduce MalAlgoQA, a real-world dataset consisting of multiple-choice questions in mathematics and reading comprehension for students in grade 3 through 11. The dataset is designed to evaluate the counterfactual reasoning abilities of LLMs through the task of analyzing the rationales and reasons behind incorrect answer choices, a process we refer to as “Malgorithm identification.” The MalAlgoQA dataset provides valuable resources to assess the counterfactual reasoning capabilities of LLMs through a pedagogical approach.
@@ 2.1 Dataset Overview
MalAlgoQA comprises two multiple-choice question sets: a mathematics and a reading comprehension question set. The mathematics question set contains 807 questions, while the reading comprehension question set has 290 questions. Each question is associated with four answer choices, as well as their corresponding rationales. Reading question set also provides a passage for each question. Table 1 provides an example of math question, with more examples available in A. Additionally, the dataset provides rich metadata including the grade level, content classification, and Depth of Knowledge (DOK) level for each question.
@@ 2.1.3 Depth of Knowledge (DOK) Levels
Each question in MalAlgoQA is annotated with a DOK level, which indicates the cognitive complexity of the question. The DOK levels are derived from Norman L. Webb’s taxonomy Webb (2002) and range from I to III, with higher levels representing increased cognitive complexity. Table 4 shows the distribution of questions across DOK levels for both question sets.
@@ 3 Methodology
Answer Rationales: {formatted_rationales}
@@
Table 5: Prompt for the Malgorithm Identification experiment, where the LLM identifies the relationship between answer choices and rationales in a multiple-choice question-answering setting. The task involves identifying the “malgorithm”, a rationale that represents flawed reasoning steps which lead to an incorrect answer choice. The prompt provides the question, an answer choice (either correct or incorrect), and all rationales, asking the LLM to identify the rationale corresponding to the given choice. The highlighted sentence instructs the model to provide step-by-step reasoning, an instruction added for the Chain-of-Thought prompting strategy.
In this section, we present the task formulation for Malgorithm Identification.
3.1 Definitions and Task Formulation
@@
We introduce the term “malgorithm” to describe a rationale underlying an incorrect answer choice. A malgorithm is composed of a series of reasoning steps, some of which are correct, while others contain errors. We refer to these erroneous steps as “mal-rules”. The term mal-rules is derived from the cognitive error literature (Brown and VanLehn, 1980; VanLehn, 1990; Payne and Squibb, 1990), which refers to the flawed or incorrect rules that students may apply when solving problems, resulting in systematic errors. In our context, a malgorithm is a reasoning chain that consists of both correct rules and mal-rules, with the presence of mal-rules ultimately leading to the selection of an incorrect answer choice.
@@ 3.2 Evaluation Metrics
1. Algorithm Identification Accuracy (AIA): This metric measures the model’s ability to identify the rationale or algorithm that corresponds to the correct answer choice. Essentially, AIA evaluates how well the model recognizes and follows a chain of reasoning steps that lead to the correct answer.
2. Malgorithm Identification Accuracy (MIA): This metric measures the model’s ability to identify the flawed reasoning or malgorithm that corresponds to an incorrect answer choice. Specifically, MIA evaluates the model’s capacity to identify faulty reasoning chains that contain a combination of correct steps and one or more mal-rules, which lead to the wrong answer.
4 Experiments and Findings
@@
Table 6: Performance results of LLMs on the Malgorithm Identification task for Reading and Math subjects. Results are organized by prompting type (Simple/CoT) and number of shots (0/5). Metrics include Multiple-Choice Question (MCQ) accuracy, Algorithm Identification Accuracy (AIA), and Malgorithm Identification Accuracy (MIA). AIA measures correct rationale identification, while MIA measures flawed reasoning identification. MIA scores are generally significantly lower than AIA, highlighting the challenge of counterfactual reasoning.
@@ 4.1 Experimental Setup
We conducted experiments on the Malgorithm Identification task with four state-of-the-art LLMs: GPT-4o, GPT-3.5, LLaMA3-70B and LLaMA3-8B.The specific models we use are GPT-4o: gpt-4o-2024-05-13, GPT-3.5: gpt-3.5-turbo-0125, LLaMA3-70B: Meta-LLaMA3-70B-Instruct, LLaMA3-8B: Meta-LLaMA3-8B-Instruct To evaluate the LLMs performance, we use AIA and MIA metrics (described in section 3.2).
@@
For each model, we evaluated with four experimental settings: (1) Simple prompting, where the LLMs are prompted to provide only the final answer; (2) Chain-of-Thought (CoT) prompting Wei et al. (2022b); Kojima et al. (2022), where the LLMs are prompted to provide step-by-step reasoning and show the intermediate steps before giving the final answer, as shown in Table 5; (3) Zero-shot learning, where the LLMs are provided without any additional examples; (4) Few-shot learning, where the LLMs are provided with a small number of examples demonstrating the desired reasoning process, in addition to the standard prompts. Due to context length constraints, we conducted few-shot experiments only on the Math dataset.
4.2 Main Findings
Wide Performance Gap Between MIA and AIA
The most striking observation from Table 6 is the substantial performance gap between Malgorithm Identification Accuracy (MIA) and Algorithm Identification Accuracy (AIA) across all models. For GPT-4o on the Math dataset with zero-shot CoT prompting, we see an AIA of $95.65\%$ compared to an MIA of only $66.10\%$ a drop of nearly $30$ percentage points. This pattern is consistent across models and prompting strategies, with the gap widening for smaller models. For instance, GPT-3.5 shows an even more dramatic difference: $86.28\%$ AIA vs $14.74\%$ MIA.
This disparity suggests that while LLMs have become proficient at identifying correct reasoning paths, they struggle significantly when asked to recognize and articulate flawed reasoning. This challenge goes beyond simple counterfactual reasoning and points to a fundamental limitation in how these models process and evaluate logical structures. The ability to identify correct reasoning does not necessarily translate to the ability to pinpoint errors in incorrect reasoning, indicating a potential blind spot in the training or architecture of current LLMs.
Impact of Model Size on MIA Performance
The data reveals a clear correlation between model size and MIA performance. For the Math dataset with zero-shot CoT prompting, we see a progression from LLaMA3-8B ($27.05\%$ MIA) to LLaMA3-70B ($55.32\%$ MIA) to GPT-4o ($66.10\%$ MIA). This trend suggests that the ability to identify malgorithms is an emergent property that significantly improves with increased model capacity Wei et al. (2022a). However, it’s important to note that this improvement with model size is much more pronounced for MIA than for AIA. While GPT-4o outperforms LLaMA3-8B by $39.05$ percentage points in MIA, the difference in AIA is only $12.58$ percentage points ($95.65\%$ vs $83.07\%$). This disproportionate scaling indicates that malgorithm identification may require more complex reasoning capabilities that are only unlocked at larger model scales.